# Bayesian inversion with Student’s $t$ priors based on Gaussian scale mixtures

### v1

@@ Abstract
Many inverse problems focus on recovering a quantity of interest that is a priori known to exhibit either discontinuous or smooth behavior. Within the Bayesian approach to inverse problems, such structural information can be encoded using Markov random field priors. Here, we propose a class of priors that combine Markov random field structure with Student’s $t$ distribution. This approach offers flexibility in modeling diverse structural behaviors depending on available data. Flexibility is achieved by including the degrees of freedom parameter of Student’s $t$ in the formulation of the Bayesian inverse problem. To facilitate posterior computations, we employ Gaussian scale mixture representation for the Student’s $t$ Markov random field prior which allows expressing the prior as a conditionally Gaussian distribution depending of auxiliary hyperparameters. Adopting this representation, we can derive most of the posterior conditional distributions in closed form and utilize the Gibbs sampler to explore the posterior. We illustrate the method with two numerical examples: signal deconvolution and image deblurring.
@@ 1 Introduction
Inverse problems aim at recovering a quantity of interest from indirect observations based on a (forward) mathematical model Tarantola (2004). Depending on the structural features of unknown to be recovered, inverse problems can be divided into different types. Some problems seek reconstruction of unknown with discontinuous structure exhibiting large jumps, while others aim at recovering unknown functions that are almost surely continuous. Examples of inverse problems requiring preservation of the edges or sharp features in the solution occur in imaging science such as X-ray computed tomography, image deblurring, segmentation and denoising Hansen (2010). Conversely, the emphasis in many inverse problems often lies in recovering smooth features, typically when determining coefficients of partial differential equations Marzouk and Najm (2009); Cui et al. (2014); Uribe et al. (2021).
@@
One of the peculiarities of inverse problems is their instability or ill-posed nature, where even minor perturbations in observations can lead to nonexistence or lack of uniqueness in the solution. To obtain a stable solution, deterministic regularization techniques such as Tikhonov regularization are often used (see, e.g., Hansen (2010)). An alternative approach reformulates inverse problems as problems of statistical inference Kaipio and Somersalo (2005). In particular, in Bayesian statistics the unknown quantity, observations and noise are treated as random variables, and the solution to the inverse problem is formulated in the form of posterior distribution.
@@
Bayesian framework naturally allows incorporating prior knowledge about the structure of unknown in the form of prior distribution. Markov random field (MRF) priors serve as a useful tool in constructing structural priors (see, e.g., Bardsley (2019)). Such models can be integrated with the Student’s $t$ distribution, offering a flexible prior capable of capturing diverse solution behaviors. The Student’s $t$ distribution can be tuned through its degrees of freedom parameter. With lower degrees of freedom, the Student’s $t$ distribution exhibits heavy-tailed characteristics (sharpness). Conversely, as the degrees of freedom parameter approaches infinity, the distribution converges towards a Gaussian distribution (smoothness).
@@
While providing enhanced flexibility, the combination of MRF priors with Student’s $t$ distributions can lead to complex posterior distributions that are hard to explore. To alleviate some of these difficulties, we utilize the Gaussian scale mixture (GSM) representation of the Student’s $t$ distribution Andrews and Mallows (1974). GSMs allow for the representation of a random variable in terms of a conditionally Gaussian variable and a mixing random variable, offering several advantages, particularly in sampling procedures. By employing the GSM for the Student’s $t$ distribution, we express it in terms of Gaussian and inverse-gamma distributions. Consequently, many conditional terms of the associated posterior simplify into products of Gaussian and inverse-gamma densities, enabling closed-form expressions through conjugacy. This construction motivates the adoption of the Gibbs sampler for posterior sampling Geman and Geman (1984).
@@
One notable challenge in our approach arises from the inference of the degrees of freedom parameter, as obtaining the conditional posterior for this parameter in closed form is unfeasible. To address this, we employ the classic Metropolis algorithm to sample the conditional within the Gibbs structure, thereby constructing a hybrid Gibbs sampler (Robert and Casella, 2004, p.389), commonly referred to as Metropolis-within-Gibbs Müller (1991). Furthermore, the statistical modeling of the degrees of freedom parameter of the Student’s $t$ distribution is a challenging task. Oftentimes this parameter is not clearly determined by the data. Finding tractable informative prior distributions for such parameter is necessary. Popular prior choices existing in statistics are: exponential prior Geweke (1993); Fernandez and Steel (1998), Jeffreys prior Fonseca et al. (2008), gamma prior Juárez and Steel (2010), penalized complexity prior Simpson et al. (2017), reference prior He et al. (2021), and log-normal prior Lee (2022). In this work, we also computationally compare some of the previous priors, aiming to offer recommendations tailored to linear inverse problem scenarios.
Related work.
@@
In linear Bayesian inverse problems, MRF priors are commonly employed to induce structural characteristics in the solution. For instance, Gaussian MRF priors are utilized to introduce smoothing structure Bardsley (2013); Roininen et al. (2014), while piecewise constant structure are modeled using Laplace MRF Bardsley (2012), Cauchy and alpha-stable MRF Suuronen et al. (2022); Senchukova et al. (2023); Suuronen et al. (2023), and horseshoe MRF Uribe et al. (2023). Additionally, hierarchical prior models, as described in Calvetti and Somersalo (2007, 2008); Calvetti et al. (2020), also adopt an MRF structure, particularly effective for representing sharp features in the solution.
@@
The modeling of the Student’s $t$ distribution in terms of a GSM has been explored in the previous literature, with seminal works dating back to Chu (1973); Andrews and Mallows (1974). Furthermore, applications of Gibbs sampling based on the GSM formulation of the Student’s $t$ have been investigated across various domains, including macroeconomic time series Geweke (1993), random effects models Choy and Smith (1997), and empirical data analysis Fernandez and Steel (1998).
Contributions and structure of the paper.
The main highlights of this work are outlined as follows
@@
    We introduce a class of flexible priors based on MRF structure combined with Student’s $t$ distribution. These priors can accommodate reconstruction of either sharp or smooth features in the solution depending on the data available.
@@
    We achieve the flexibility in modeling different structural behaviour by solving a hierarchical Bayesian inverse problem that also estimates the degrees of freedom parameter of the Student’s $t$ distribution, for which we test different types of informative priors.
@@
    We exploit GSM representations of Student’s $t$ to sample the resulting posterior more efficiently using the Gibbs sampler.
@@
    We compare the accuracy and performance of the Gibbs sampler with the results obtained by Hamiltonian Monte Carlo based on its No-U-Turn Sampler (NUTS) adaptation Hoffman and Gelman (2014). We demonstrate the performance of our method through numerical experiments on linear inverse problems.
@@ 2.1 Discrete Bayesian inverse problems
We consider the inverse problem of finding unknown parameter functions of a mathematical model that are defined over a domain $\mathcal{D}$, using noisy observed data of an underlying physical process that the mathematical model describes. This represents an infinite-dimensional inverse problem that we approach by first discretizing all functions and then using a finite-dimensional perspective. In this case, the dimension $d$ of the discretized function will depend on the discretization size $N$, for example, in one- and two-dimensional problems $d=N$ and $d=N^{2}$ (assuming equal discretization in both directions), respectively.
We employ the Bayesian approach to inverse problems Kaipio and Somersalo (2005). The unknown parameter function is represented as a discretized random field given by a random vector $\bm{X}$ taking values $\bm{x}\in\mathcal{X}:=\mathbb{R}^{d}$ indexed on $\mathcal{D}$. Hence, we assume the distribution of $\bm{X}$ is absolutely continuous with respect to the Lebesgue measure, and it has a well-defined prior probability density $p(\bm{x})$.
@@
where the noise covariance matrix is defined as $\bm{\Sigma}_{\mathrm{obs}}:=\sigma_{\mathrm{obs}}^{2}\bm{I}_{m}$ with the noise variance $\sigma_{\mathrm{obs}}^{2}$, and $\bm{I}_{m}$ is the identity matrix of size $m$. From the Gaussian model used for the noise and further assuming that the noise and model parameters are independent, the likelihood is defined as
$p(\bm{y}\;|\;\bm{x})=\frac{1}{(2\pi)^{\nicefrac{{m}}{{2}}}\sigma_{\mathrm{obs}}^{m}}\exp\left(-\frac{1}{2\sigma_{\mathrm{obs}}^{2}}\left\|\bm{y}-\bm{A}(\bm{x})\right\|_{2}^{2}\right).$ (2)
@@
Having defined the posterior, we are typically faced with the task of extracting information from it. This is in general not a trivial task in high dimensions. One practical option is to employ variational methods, which assume that most of the information resides in a single peak of the posterior, e.g., the maximum a posteriori (MAP) estimator, which is a point $\bm{x}_{\scriptscriptstyle\rm MAP}$ that maximizes eq. 3. Depending on the forward operator and the choice of the prior, this task can be a non-convex optimization problem; we likely find local maximizers that can still be used to build approximations to the posterior. Another method for characterizing the posterior in high dimensions is sampling. The idea is to generate a set of points $\{\bm{x}_{i}\}_{i=1}^{n}$ distributed, sometimes only approximately, according to $p\left(\bm{x}\;|\;\bm{y}\right)$. This enables the estimation of several summary statistics via Monte Carlo methods, in particular the posterior mean (PM), which is an optimal point estimate $\bm{x}_{\scriptscriptstyle\rm PM}$ in the sense that it minimizes the Bayes risk (see, e.g., (Tenorio, 2017, p.88)).
Our focus is to estimate posterior distributions arising in imaging science, including cases where the unknown solutions can have either sharp or smooth features. These characteristics are specified via prior distributions that have the flexibility to express different behaviour of the unknown.
@@ 2.2 Encoding structural information
In infinite-dimensional Bayesian inverse problem, priors are usually constructed directly in functional spaces. Following this approach, Gaussian priors promoting the smooth behavior can be represented using, e.g., the Karhunen-Loève expansion Spanos and Ghanem (1989). Besov space priors are defined on a wavelet basis and, unlike Gaussian priors, are able to produce discontinuous samples, which is useful in the problems requiring edge preservation Lassas et al. (2009). Neural networks can also be used to impose smooth or sharp behaviors depending on the distribution chosen for the weights and biases of the network Li et al. (2022).
Since our focus is on discrete Bayesian inverse problems, we follow a different approach in which the functional space is first discretized and then priors are constructed. Under this formulation, MRF models can be used to incorporate structural information into the prior. MRFs are defined in connection with an undirected graph that contains vertices and edges associating dependencies between elements of the solution vector. For example, when the distribution of the MRF is Gaussian, the graph is defined by the nonzero pattern of the precision matrix (Rue and Held, 2005, p.21). When there is strong dependency between neighbouring elements of the solution, a common way to define the precision matrix is via the forward difference matrix (with zero boundary condition)
@@
where $N$ is the discretization size. In two-dimensional applications, we often distinguish between horizontal and vertical differences, $\bm{D}^{(1)}=\bm{I}_{N}\otimes\bm{D},\quad\bm{D}^{(2)}={\bm{D}}\otimes\bm{I}_{N}$, respectively ($\otimes$ denotes the Kronecker product and $\bm{I}_{N}$ is a square identity matrix of size $N$). We remark that in practice one works with intrinsic Gaussian MRFs that have precision matrices not of full rank. Moreover, since we utilize first-order forward differences, the intrinsic MRF is known to be of first order as its precision matrix has rank $d-1$ (Rue and Held, 2005, p.87).
Using the pairwise difference structure set via eq. 4, we can define a new parameter vector $\bm{u}$ containing differences
@@
Then the posterior eq. 3 can be reformulated as one of the following densities
$p\left(\bm{u}\;|\;\bm{y}\right)$ $\propto p(\bm{y}\;|\;\bm{L}^{-1}\bm{u},{\lambda})\,p(\bm{u}),$ (6a)
$p\left(\bm{x}\;|\;\bm{y}\right)$ $\propto p(\bm{y}\;|\;\bm{x},{\lambda})\,p(\bm{L}\bm{x}).$ (6b)
The posterior eq. 6a uses a reparameterization of the inverse problem in terms of the unknown parameter differences $\bm{u}$ which facilitates computations involving the prior term, however, it does complicate the likelihood term as one requires inverting the difference matrix $\bm{L}$, which is not straightforward in other than one-dimensional problems (see, e.g., Calvetti and Somersalo (2007)). The posterior eq. 6b keeps the inverse problems in terms of the unknown parameter $\bm{x}$, however, computations of the prior term are more involved (see, e.g.,Bardsley (2019)). Both densities are targeting the same problem, however, we focus on the second parameterization as it does not require especial methods to invert matrix $\bm{L}$.
We have seen that MRF models can be used to set a prior that incorporates strong dependence information. We still require the specification of the probability distribution of the MRF.
Since we aim at defining a flexible model that can also account for smooth features in the solution, we consider a special yet rather wide class of distributions — Student’s $t$ distribution. Our intention is to combine Student’s $t$ with the MRF structure presented in this subsection. The main advantage is that such distribution family can be controlled by a degrees of freedom parameters defining how heavy-tailed the distribution is. Samples from heavy-tailed distributions have larger jumps, compared to those sampled from Gaussian. The main difficulty is that resulting prior models can complicate the characterization of high-dimensional posteriors. To circumvent this, we explore GSM representations, which are discussed next.
2.3 Gaussian scale mixture representation
The posterior can be examined in some depth by exploiting certain properties of the assumed priors. This is the case when the prior distributions can be constructed as GSM Chu (1973); Andrews and Mallows (1974). This class of distributions has densities that are expressed as an integral of a Gaussian density. The random variable being expressed as a GSM is usually referred to as sub-Gaussian, i.e., its distribution has strong tail decay, at least as fast as the tails of a Gaussian (Samorodnitsky and Taqqu, 1994, p.77).
Let $Y\sim\mathcal{N}(0,1)$ be a standard Gaussian random variable, and $Z$ a positive random variable with density $p_{\scriptscriptstyle Z}(z)$. Moreover, $Y$ and $Z$ are independent. We say that the density of the random variable $X=Y\,Z^{\nicefrac{{1}}{{2}}}$ is a GSM with mixing density $p_{\scriptscriptstyle Z}(z)$, and it is given by

### v2

@@ Abstract
Many inverse problems focus on recovering a quantity of interest that is a priori known to exhibit either discontinuous or smooth behavior. Within the Bayesian approach to inverse problems, such structural information can be encoded using Markov random field priors. We propose a class of priors that combine Markov random field structure with Student’s $t$ distribution. This approach offers flexibility in modeling diverse structural behaviors depending on available data. Flexibility is achieved by including the degrees of freedom parameter of Student’s $t$ distribution in the formulation of the Bayesian inverse problem. To facilitate posterior computations, we employ Gaussian scale mixture representation for the Student’s $t$ Markov random field prior, which allows expressing the prior as a conditionally Gaussian distribution depending on auxiliary hyperparameters. Adopting this representation, we can derive most of the posterior conditional distributions in a closed form and utilize the Gibbs sampler to explore the posterior. We illustrate the method with two numerical examples: signal deconvolution and image deblurring.
@@ 1 Introduction
Inverse problems aim at recovering a quantity of interest from indirect observations based on a (forward) mathematical model [1]. Depending on the structural features of the unknown to be recovered, inverse problems can be divided into different types. Some problems seek reconstruction of the unknown with discontinuous structures exhibiting large jumps, while others aim at recovering unknown functions that are almost surely continuous. Examples of inverse problems requiring preservation of the edges or sharp features in the solution occur in imaging science, such as X-ray computed tomography, image deblurring, segmentation, and denoising [2, 3, 4]. Conversely, the emphasis in many inverse problems often lies in recovering smooth features, typically when determining coefficients of partial differential equations [5, 6, 7].
@@
One of the characteristics of inverse problems is their instability or ill-posed nature, where even minor perturbations in observations can lead to nonexistence or lack of uniqueness in the solution. To obtain a stable solution, deterministic regularization techniques such as Tikhonov regularization are often used (see, e.g., [4]). An alternative approach reformulates inverse problems as problems of statistical inference [8]. In particular, in Bayesian statistics, the unknown quantity, observations, and noise are treated as random variables, and the solution to the inverse problem is formulated in the form of the posterior distribution.
@@
Bayesian framework naturally allows incorporating prior knowledge about the structure of the unknown, which also acts as a regularizer, in the form of the prior distribution. Markov random field (MRF) priors are a useful tool in constructing structural priors (see, e.g., [9]). Such models can be integrated with Student’s $t$ distribution, offering a flexible prior capable of capturing diverse solution behaviors. Student’s $t$ distribution can be tuned through its degrees of freedom parameter. With lower degrees of freedom, Student’s $t$ distribution exhibits heavy-tailed characteristics (sharpness). Conversely, as the degrees of freedom parameter approaches infinity, the distribution converges towards a Gaussian distribution (smoothness).
@@
While providing enhanced flexibility, the combination of MRF priors with Student’s $t$ distributions can lead to complex posterior distributions that are hard to explore. To alleviate some of these difficulties, we utilize the Gaussian scale mixture (GSM) representation of Student’s $t$ distribution [10]. GSMs allow for the representation of a random variable in terms of a conditionally Gaussian variable and a mixing random variable, offering several advantages, particularly in sampling procedures to explore the posterior. By employing the GSM for Student’s $t$ distribution, we can express it in terms of conditionally Gaussian and inverse-gamma distributions. Consequently, many conditional terms of the resulting hierarchical posterior simplify into products of Gaussian and inverse-gamma densities, enabling closed-form expressions through conjugacy. This construction motivates the adoption of the Gibbs sampler for posterior sampling [11].
@@
One notable challenge in our approach arises from the inference of the degrees of freedom parameter, as obtaining the conditional posterior for this parameter in closed form is unfeasible. To address this, we employ the classic Metropolis algorithm to sample the conditional within the Gibbs structure, thereby constructing a hybrid Gibbs sampler [12, p.389], commonly referred to as Metropolis-within-Gibbs [13]. Furthermore, the statistical modeling of the degrees of freedom parameter of Student’s $t$ distribution is a challenging task. Oftentimes this parameter is not clearly determined by the data. Finding tractable informative prior distributions for such a parameter is necessary. Popular prior choices existing in statistics are exponential prior [14, 15], Jeffreys prior [16], gamma prior [17], penalized complexity prior [18], reference prior [19], and log-normal prior [20]. In this work, we also computationally compare some of the previous priors, aiming to offer recommendations tailored to linear inverse problem scenarios.
Related work
@@
In linear Bayesian inverse problems, MRF priors are commonly employed to induce structural characteristics in the solution. For instance, Gaussian MRF priors are utilized to introduce smoothing structure [21, 22], while piecewise constant structure are modeled using Laplace MRF [23], Cauchy and alpha-stable MRF [24, 25, 26], and horseshoe MRF [27]. Additionally, hierarchical prior models, as described in [28, 29, 30], also adopt an MRF structure, particularly effective for representing sharp features in the solution.
@@
The modeling of Student’s $t$ distribution in terms of a GSM has been explored in the previous literature, with seminal works dating back to [31, 10]. So-called hierarchical $t$ formulations for MRFs are briefly discussed in [32, Ch. 5]. Furthermore, applications of Gibbs sampling based on the GSM formulation of Student’s $t$ distribution have been investigated across various domains, including macroeconomic time series [14], random effects models [33], and empirical data analysis [15].
Contributions and structure of the paper
The main highlights of this work are outlined as follows:
@@
    We introduce a class of flexible priors based on MRF structure combined with Student’s $t$ distribution. These priors can accommodate the reconstruction of either sharp or smooth features in the solution depending on the available data.
@@
    We reach such flexibility by solving a hierarchical Bayesian inverse problem that also estimates the degrees of freedom parameter of Student’s $t$ distribution, for which we analyze different types of informative priors.
@@
    We exploit the GSM representation of Student’s $t$ distribution to sample the resulting hierarchical posterior more efficiently using the Gibbs sampler.
@@
    We compare the accuracy and performance of the Gibbs sampler with the results obtained by Hamiltonian Monte Carlo based on its No-U-Turn Sampler (NUTS) adaptation [34]. The performance of our method is demonstrated through numerical experiments on linear inverse problems.
@@ 2.1 Discrete Bayesian inverse problems
We consider the inverse problem of finding unknown parameter functions of a mathematical model that are defined over a domain $\mathcal{D}$, using noisy observed data of an underlying physical process described by the mathematical model. This represents an infinite-dimensional inverse problem that we approach by first discretizing all functions and then using a finite-dimensional perspective. In this case, the dimension $d$ of the discretized function will depend on the discretization size $N$, for example, in one- and two-dimensional problems $d=N$ and $d=N^{2}$ (assuming equal discretization in both directions), respectively.
We employ the Bayesian approach to inverse problems [8]. The unknown parameter function is represented as a discretized random field given by a random vector $\bm{X}$ taking values $\bm{x}\in\mathcal{X}:=\mathbb{R}^{d}$ indexed on $\mathcal{D}$. Hence, we assume the distribution of $\bm{X}$ is absolutely continuous with respect to the Lebesgue measure, and it has a well-defined prior probability density $p(\bm{x})$.
@@
where the noise covariance matrix is defined as $\bm{\Sigma}_{\mathrm{obs}}:=\sigma_{\mathrm{obs}}^{2}\bm{I}_{m}$ with noise variance $\sigma_{\mathrm{obs}}^{2}$ and identity matrix $\bm{I}_{m}$ of size $m$. Since the noise is assumed to be Gaussian, and further assuming that the noise and model parameters are independent, the likelihood is defined as
$p(\bm{y}\;|\;\bm{x})=\frac{1}{(2\pi)^{\nicefrac{{m}}{{2}}}\sigma_{\mathrm{obs}}^{m}}\exp\left(-\frac{1}{2\sigma_{\rm obs}^{2}}\left\|\bm{y}-\bm{A}(\bm{x})\right\|_{2}^{2}\right).$ (2)
@@
Having defined the posterior, we are typically faced with the task of extracting information from it. This is in general not a trivial task in high dimensions. One practical option is to employ variational methods, which assume that most of the information resides in a single peak of the posterior, e.g., the maximum a posteriori (MAP) estimator, which is a point $\bm{x}_{\scriptscriptstyle\rm MAP}$ that maximizes posterior eq. 3. Depending on the forward operator and the choice of the prior, this task can be a non-convex optimization problem; we likely find local maximizers that can still be used to build approximations to the posterior. Another method for characterizing the posterior in high dimensions is sampling. The idea is to generate a set of points $\{\bm{x}_{i}\}_{i=1}^{n}$ distributed, sometimes only approximately, according to $p\left(\bm{x}\;|\;\bm{y}\right)$. This enables the estimation of several summary statistics via Monte Carlo methods, in particular the posterior mean (PM), which is an optimal point estimate $\bm{x}_{\scriptscriptstyle\rm PM}$ in the sense that it minimizes the Bayes risk (see, e.g., [35, p.88]).
Our focus is to estimate posterior distributions arising in imaging science, including cases where the unknown solutions can have either sharp or smooth features. These characteristics can be modeled via prior distributions that have the flexibility to express different features in the unknown solution.
Remark 2.1.
The interplay between the noise variance and the (scale-squared) parameters of the prior essentially defines the so-called regularization parameter used in classical deterministic inverse problems to control the quality of the solution. In practice, however, the parameter $\sigma_{\rm obs}^{2}$ is usually unknown, and hence, we decide to model it as a random variable and infer it from the data. In this case, we use an inverse-gamma hyperprior $\sigma_{\rm obs}^{2}\sim\text{IG}(a,b)$, with shape parameter $a=1$ and scale parameter $b=10^{-4}$, following the recommendation in [36].
@@ 2.2 Encoding structural information
In infinite-dimensional Bayesian inverse problems, priors are usually constructed directly in functional spaces. Following this approach, Gaussian priors promoting the smooth behavior can be represented using, e.g., the Karhunen–Loève expansion [37]. Besov space priors are defined on a wavelet basis and, unlike Gaussian priors, are able to produce discontinuous samples, which is useful in problems requiring edge preservation [38]. Neural networks can also be used to impose smooth or sharp behaviors depending on the distribution chosen for the weights and biases of the network [39].
Since our focus lies in discrete Bayesian inverse problems, we follow a different approach in which the involved functions are first discretized, and then priors are constructed on those discretizations. Under this formulation, MRF models can be used to incorporate structural information into the prior. MRFs are defined in connection with an undirected graph that contains vertices and edges associating dependencies between elements of the solution vector. For example, when the distribution of the MRF is Gaussian, the graph is defined by the nonzero pattern of the precision matrix [32, p.21]. When there is a strong dependency between neighboring elements of the solution, a common way to define the precision matrix is via the forward difference matrix (with zero boundary condition)
@@
where $N$ is the discretization size. In two-dimensional applications, we often distinguish between horizontal and vertical differences, $\bm{D}^{(1)}=\bm{I}_{N}\otimes\bm{D},\quad\bm{D}^{(2)}={\bm{D}}\otimes\bm{I}_{N}$, respectively ($\otimes$ denotes Kronecker product and $\bm{I}_{N}$ is a square identity matrix of size $N$). We remark that in practice one works with intrinsic Gaussian MRFs, which have precision matrices not of full rank. Moreover, since we utilize first-order forward differences, the intrinsic MRF is known to be of the first order as its precision matrix has rank $d-1$ [32, p.87].
Using the pairwise difference structure set via Equation eq. 4, we can define a new parameter vector $\bm{u}$ containing differences
@@
Then, the posterior eq. 3 can be reformulated as one of the following densities
$p\left(\bm{u}\;|\;\bm{y}\right)$ $\propto p(\bm{y}\;|\;\bm{L}^{-1}\bm{u},\sigma_{\rm obs}^{2})\,p(\bm{u}),$ (6a)
$p\left(\bm{x}\;|\;\bm{y}\right)$ $\propto p(\bm{y}\;|\;\bm{x},\sigma_{\rm obs}^{2})\,p(\bm{L}\bm{x}).$ (6b)
The posterior eq. 6a uses a reparameterization of the inverse problem in terms of the unknown parameter differences $\bm{u}$, which facilitates computations involving the prior term; however, it does complicate the likelihood term as one requires inverting the difference matrix $\bm{L}$, which is not straightforward in other than one-dimensional problems (see, e.g., [28]). The posterior eq. 6b keeps the unknown parameter $\bm{x}$ as an objective of inference; however, computations of the prior term are more involved (see, e.g.,[9]). Both formulations are essentially targeting the same problem, however, we focus on the second parameterization, as it does not require special methods to invert matrix $\bm{L}$.
We have seen that MRF models can be used to set a prior that incorporates strong dependence information. We still require the specification of the probability distribution of the MRF. Since we aim at defining a flexible model that can also account for smooth features in the solution, we consider a special yet rather wide class of distributions — the Student’s $t$ distribution. We intend to combine Student’s $t$ distribution with the MRF structure presented in this subsection. The main advantage is that such a distribution family can be controlled by the degrees of freedom parameter defining how heavy-tailed the distribution is. Samples from heavy-tailed distributions have larger jumps, compared to those sampled from Gaussian. The main difficulty is that resulting prior models can complicate the characterization of high-dimensional posteriors. To circumvent this, we explore GSM representations, which are discussed next.
2.3 Gaussian scale mixtures
The posterior can be examined in more depth by exploiting certain properties of the assumed priors. This is the case when the prior distributions can be expressed as a Gaussian scale mixture, also known as scale mixtures of normals [40, 31, 10]. The GSM formulation plays a major role in hierarchical modeling, as it allows the expression of more involved distributions in a hierarchical manner by taking an integral of a Gaussian density.
Definition 1 (Gaussian scale mixture).
Let $Y\sim\mathcal{N}(0,1)$ be a standard Gaussian random variable, and $Z$ a positive random variable with density $p_{\scriptscriptstyle Z}(z)$. Moreover, $Y$ and $Z$ are independent. We say that the density of the random variable $X=Y\,Z^{\nicefrac{{1}}{{2}}}$ given by